# Revisiting Point Cloud Completion: Are We Ready For The Real-World?

### v2

@@ 4.1 Comparative Analysis
In this section, we conduct a comprehensive geometric and topological analysis and comparison of our dataset with PCN and ShapeNet-55/34 both qualitatively and quantitatively based on the following three factors:
  • Non-Uniformity: For a point cloud we calculate the distance of all of its N points with their nearest neighbor and then find the standard deviation of all these distances across the entire point cloud. We report the average of all these standard deviation values across different classes of different datasets as Non-Uniformity.
  • Noise: For each point in a point cloud we first fit a plane using linear regression over its k-nearest neighbors. Then we calculate the perpendicular distance from the said point to the said plane. We average this distance for all the points in our point cloud. Finally, we calculate the average of these average values, termed Noise, across different classes of different datasets.
@@
  • Persistent Homology ($H_{0}$; $H_{1}$): Persistence Diagrams in $\mathcal{PH}$ represent the $(birth,death)$ of topological features across a filtration, as points on the x-y plane. All points are above the diagonal $(x=y)$. Points further away from the diagonal represent features with a larger persistence ($(death-birth)$), and hence are the most important topological features. We compute the average of $death-birth$ averages for 0 & 1-dim topological features separately across all classes of all datasets and report them as PH-based ($H_{0}$; $H_{1}$).
@@
We pick four classes at random from PCN and ShapeNet datasets. Whereas, we divide our RealPC into four classes based on their sources. As presented in Table 1, our dataset shows the highest Noise. Non-uniformity, and PH-based ($H_{0}$; $H_{1}$) values consistently when compared to any of the other synthetic classes. This result is supported by the visualizations shown in Figure 1 (a) and (b). We can clearly see high non-uniformity and noise in point clouds from our dataset, in contrast to the complete absence of them in PCN. The (b) part of this figure investigates the existence of topological features in real and synthetic point clouds. The contrast between the two is striking. While persistence diagram for ours consists of numerous important 0 & 1-dim topological features, synthetic datasets (e.g. PCN) lack them. This is indicated by the presence of several points further from the diagonal for our dataset and their absence in PCN.
5 Experiments
Point Cloud completion is one of the fundamental tasks in point clouds perception-based challenges. In this section, we demonstrate the challenges in real-world point completion by benchmarking RealPC against several strong baselines. In order to gauge the maturity of the point cloud completion literature we benchmark them against several recent, strong, and relevant baselines that have shown remarkable performance on existing simulated datasets. We demonstrate that current state-of-the-art models are not adept at handling point clouds collected in real-world environments that have several contrasting characteristics from real-world point clouds. We demonstrate the performance of these models in Table 2.
5.1 Benchmarking
We benchmark baselines that have shown remarkable performance on existing datasets (ShapeNet, PCN, ShapeNet33, ShapeNet55). We use the following works and show their performance on RealPC: (a) ODGNet [6] (b) PointAttN [49](c) SnowFlakeNet [52] (d) PointTr [57] (e) AdaPoinTr [57](f) PartialToComplete [16](g) TopNet [44] (h) GrNet [54] (h) SeedFormer[63]. We use Chamfer Distance (L1) as a metric to measure the conformity of the completed shapes against the ground truth complete shapes.
5.2 Results
We now demonstrate the results of our benchmarking exercise in Table 2 and Figure 5. We demonstrate the completion performance of all the available baselines on RealPC across the available 21 classes. The results across all models demonstrate the challenges presented by RealPC. The benchmarked models fail miserably while generating complete shapes of partial incomplete point clouds from the RealPC dataset. These results demonstrate the gap between performance on real-world point clouds and current simulated point clouds that have been used in the literature.
5.3 Analysis
Refer to caption Refer to captionRefer to caption
Figure 5: Left: Non-uniform sparse partial input to the best benchmarked baseline. Middle: Sub-optimal completion results. Right: Ground Truth complete point cloud.
@@ 5.3 Analysis
We discuss the performance of the benchmarked baseline models on RealPC dataset. As demonstrated in Table 2 and Figure 5, we see that the current best models are in-adept at precise point cloud completion. We discuss the reasons for this behavior - (1) We investigate synthetic and real-world world point clouds using tools from algebraic topology and topological data analysis in Section 4.1. Our findings, though striking are not surprising. The observation suggests that RealPC consists of non-trivial 0 and 1-dimensional Homology-based topological features that correspond to connected components, and cycles. This is indicated by the non-diagonal nature of the persistence diagram in Figure 1. These topological features add to the complexity of RealPC and provide strong evidence of the existence of higher dimensional topological features. We demonstrate through experiments that adding these Homology priors as constraints to an existing model can improve completion to some degree (Sec. 6). (2) Point cloud in RealPC are collected from several different sensors. RealPC utilizes four different sensors for collecting the point clouds. This is a realistic setting in the real world where different sources of the incomplete point clouds may use sensors with different intrinsic characteristics. Different point cloud samples in such a case may follow different distributions - which is not true in the synthetic datasets. (3) We demonstrate in Section 4.1 that real-world point clouds have several characteristics that make them challenging to work with compared to synthetic point clouds. These characteristics - (non-uniformity, noise, topological features) have a relatively uniform pattern in synthetic point clouds (Table 2). On the contrary, a real-world point cloud acquisition pipeline is affected by several ungovernable parameters. These factors introduce non-uniform and inconsistent patterns that are challenging to model.
Baseline RealPC Categories
Ch0 Ch1 Ch2 Ch3 Du0 Du1 Du2 Du3 Du4 Hu0 Hu1 Hu2 Hu3 Hu4 Hu5 Hu6 Hu7 Sn0 Sn1 Sn2 Sn3 Mean
ODG 96 91 90 100 170 172 208 154 191 89 114 106 120 116 120 134 119 83 67 79 80 112
PoinTr 110 113 101 104 99 113 98 92 105 191 146 165 106 132 194 94 121 117 151 144 151 114
AdaPoinTr 65 71 71 68 57 54 52 98 37 158 132 114 75 79 134 67 86 65 47 61 53 69
FoldingNet 158 163 181 246 159 143 203 259 161 181 205 154 221 183 235 219 211 142 107 102 74 167
PCN 145 136 132 123 158 146 167 181 118 170 129 150 139 126 196 118 151 148 106 125 113 143
TopNet 563 385 200 211 88 77 90 91 45 590 110 328 94 79 441 93 93 455 75 320 129 341
SnowFlakenet 58 - 49 54 73 57 - - - - - 68 - - - - - 57 - 82 60
GRNet 79 80 79 104 83 63 66 111 64 140 133 131 92 74 155 74 82 81 71 94 73 84
Anchorformer 64 68 75 115 58 51 62 84 36 146 131 108 79 85 151 70 92 78 79 75 91 87
Table 2: Benchmark results on the RealPC. Despite remarkable performance on synthetic datasets, the baselines fail on RealPC
6 Persistent Homology Regularized Completion
We establish in Section 4.1 that RealPC consists of rich topological features while synthetic features lack these.
In this section, we focus on validating whether using topological priors on a suitable benchmarked work can improve the completion performance.
A challenge with using $\mathcal{PH}$ is that it is compute-intensive. The computation of persistence and simplicies in the simplicial complex generated using the point clouds increases exponentially with the number of points. This makes it difficult to integrate with large-scale point cloud applications. We select a Point Cloud completion model that can be easily integrated with $\mathbf{PH}$. We choose ODGNet [6] as our reference model to integrate $\mathcal{PH}$. We provide our reasons for selecting ODGNet as the reference model - ODGNet is an encoder-decoder network that works by predicting multi-scale features at the decoder. These are used to extract seed point clouds at different resolutions. ODGNet works by progressively up-scaling these at the decoder, and finally predicting the original point cloud. The sparse seed points clouds can be easily integrated with simple topological priors with moderate compute requirements. The availability of such point clouds as learnable parameters at intermediate layers in ODGNet also helps in training using the $\mathcal{PH}$ based topological loss function.
6.1 Architecture
@@
We describe the methodology for the integration of ODGNet with topological priors -TOPODGnet . We extract the multi-level seeds from the mid-level, low-level, and global features generated at the decoder. Seed points clouds consisting of points in the range of 256 to 1024 can be extracted at the decoder (Figure 6). The coarse nature of these seeds makes them easy to work with, as well as allows efficient computation of topological features.
@@
Figure 6: Architecture diagram of TOPODGnet. We calculate 0-dim $\mathcal{PH}$ based topological priors over the sparse seeds and integrate it into the loss function. It enables completion along a topologically consistent skeleton.
@@ 6.1.1 Topological Loss
Figure 7: Left: A complete point cloud Middle: Visualization of the 00-dim $\mathcal{PH}$ based skeleton on a sparse seed during initial training. Right: .00-dim $\mathcal{PH}$ skeleton using sparse seeds as the training progresses. These priors assist the model in generating topologically consistent complete point clouds.
Given an incomplete point cloud, our goal is to augment new points that follow the global topology of the ground truth complete point cloud. We extract the coarse point cloud seeds at the decoder. Our aim is to ensure that the completed point cloud follows the global topology of the ground truth complete point cloud available in the dataset. Part of this work is already handled by the dictionary module of ODGNet[6]. It ensures that the seed point clouds are able to capture the artifacts of the complete point cloud.
@@
We now apply topological regularization priors on the seed point clouds. To obtain this global topology-based backbone we use 0-dim $\mathcal{PH}$ priors on the seed point clouds. 0-dim $\mathcal{PH}$ ensures the extraction of a complete point cloud skeleton (Fig. 7) that can serve as a prior to the network to follow, for the complete point cloud generation. The underlying idea is to ensure the shape consistency of the complete output along the skeleton by generating points along the skeleton. We explain the process that extracts 0-dim $\mathcal{PH}$ priors from the seeds - Given a seed point cloud we convert the seeds into a simplicial complex. The initial complex is just the original point clouds with $\alpha$-radius ball where $\alpha$=0. Increasing $\alpha$ adds edges, faces to the simplical complex, thereby leading to the evolution and death of k-dim topological features. Each k-dim topological feature is characterized by a (birth,death) pair.
@@
Our goal is to use 0-dim $\mathcal{PH}$ based topological features. These focus on the global topology of the point cloud, thereby providing a global skeleton that outlines the complete point cloud (Figure 7).
@@
Each of the 0-dim homology features generated using filtration consists of a separate set of (birth,death) pairs denoted as b,d. The persistence of each pair is (b-d).
@@
Now we explain our topological loss function - For 0-dim $\mathcal{PH}$, we minimize the sum of persistence values for the pairs. This ensures that at the end of the filtration, there is one connected component left (given that we minimize the persistence of all the pairs and there always exists at least 1 component by default in the persistence diagram). This persistence diagram outlines the skeleton of the complete point cloud, which when used as a regularizer to a model, guides point generation along the skeleton. Given $(b_{i},d_{i})$ are the $(birth,death)$ pairs of a given seed point cloud, the topology loss ${L}_{homo}$ is defined as follows -
@@
We explain the significance of k now. The input partial point cloud may not necessarily have a single component- but may be split into more than one component (Fig. 6). In such cases, using k $\geq$ 2, allows $\mathcal{PH}$ to generate multiple skeletons - each of which can attend to partial components in the input.
6.2 Results

### v3

@@ 4.2 Comparative Analysis
We conduct a comprehensive geometric and topological analysis and comparison of RealPC with PCN and ShapeNet based on the following three factors:
@@
  • Persistent Homology ($H_{0}$; $H_{1}$): Persistence Diagrams in $\mathcal{PH}$ represent the $(birth,death)$ of topological features across a filtration, as points on the x-y plane. All points are above the diagonal $(x=y)$. Points further from the diagonal represent features with larger persistence ($death-birth$), and hence are more important. We compute the average of persistence averages, for 0- and 1-dimensional topological features, across all classes of a given dataset and report them as PH-based ($H_{0}$; $H_{1}$).
  • Non-Uniformity: For a PC we calculate the distance of all of its N points with their nearest neighbor and then find the standard deviation of all these distances across the entire PC. We report the average of all these standard deviation values across different classes of different datasets as Non-Uniformity.
  • Noise: For each point in a PC we first fit a plane using linear regression over its k-nearest neighbors. Then we calculate the perpendicular distance from the said point to the said plane. We average this distance for all the points in our PC. Finally, we calculate the average of these average values, termed Noise, across different classes of different datasets.
@@
For all three mentioned datasets, we present in Table 1 the above three metrics over four random classes and also the mean of all the classes (not just four). Our dataset shows the highest Noise, Non-uniformity, and PH-based ($H_{0}$; $H_{1}$) values consistently when compared to existing datasets. This result is supported by the visualizations shown in Figure 1 (a) and (b). We can clearly see high non-uniformity and noise in PCs from our dataset, in contrast to the complete absence of them in PCN. The (b) part of this figure investigates the existence of topological features in RealPC and existing PCs. The contrast between the two is striking. While the persistence diagram for ours consists of numerous important 0- and 1-dimensional topological features, existing datasets (e.g. PCN) lack them. This is indicated by the presence of several points further away from the diagonal for RealPC, and their absence for PCN.
4.3 Benchmarking on Non-Neural Methods
[…]
As shown in Table 2, RealPC exhibits significantly higher errors when compared to ShapeNet for the three tasks, demonstrating the complexity of RealPC. For a visual comparison, please refer to the supplementary material.
Simplification Surface Reconstruction Upsampling
CD HD CD
ShapeNet 0.522 68.198 0.043
RealPC 8.817 494.591 0.838
Table 2: Comparison of RealPC with ShapeNet on three popular non-neural tasks. We compare simplified/meshed/upsampled PCs with original PCs averaged across all classes. CD is Chamfer Distance and HD is Hausdorff Distance. All numbers are (×$\times 10^{-3}$).
ShapeNet RealPC
CD 67 90
EMD 61 90
ShapeNet RealPC
FoldingNet 5 27
DGCNN 49 58
SnowflakeNet 19 23
Table 3: Left Table: Benchmarking result for shape generation on diffusion-guided shape generation [50] using the 1-NNA metric (with both Chamfer distance (CD) and Earth Mover Distance (EMD)) as our main metric. It quantifies the distributional similarity between generated shapes and validation sets and measures both quality and diversity. Right Table: Benchmarking result for reconstruction using standard PC backbones using the Chamfer Distance metric. Both tables use Chamfer Distance (×$\times 10^{-3}$).
4.4 Benchmarking on Neural Methods
RealPC PCN
Ch0 Ch1 Ch2 Ch3 Du0 Du1 Du2 Du3 Du4 Hu0 Hu1 Hu2 Hu3 Hu4 Hu5 Hu6 Hu7 Sn0 Sn1 Sn2 Sn3 Mean (L1) Mean (L2) Mean (L1)
ODGNet 96 91 90 100 170 172 208 154 191 89 114 106 120 116 120 134 119 83 67 79 80 119 111 6
PoinTr 110 113 101 104 99 113 98 92 105 191 146 165 106 132 194 94 121 117 151 144 151 114 58 8
AdaPoinTr 65 71 71 68 57 54 52 98 37 158 132 114 75 79 134 67 86 65 47 61 53 69 26 7
FoldingNet 158 163 181 246 159 143 203 259 161 181 205 154 221 183 235 219 211 142 107 102 74 167 127 14
PCN 145 136 132 123 158 146 167 181 118 170 129 150 139 126 196 118 151 148 106 125 113 143 92 10
TopNet 563 385 200 211 88 77 90 91 45 590 110 328 94 79 441 93 93 455 75 320 129 341 1103 12
SnowflakeNet 58 - 49 - 54 73 57 - - - - - 68 - - - - - 57 - 82 60 72 7
GRNet 79 80 79 104 83 63 66 111 64 140 133 131 92 74 155 74 82 81 71 94 73 84 27 9
AnchorFormer 64 68 75 115 58 51 62 84 36 146 131 108 79 85 151 70 92 78 79 75 91 72 28 7
Table 4: Performance of baselines on RealPC dataset. We also show the corresponding average values for the PCN dataset. Despite remarkable performance on current existing datasets, the baselines fail on RealPC. We use Chamfer distance (×$\times 10^{-3}$).
In this subsection, we study and compare RealPC’s performance against existing datasets on three neural methods: (a) Completion, (b) Reconstruction, and (c) Generation. PC completion is a fundamental task in PC perception-based challenges, and it will serve as our primary focus. We benchmark several completion methods on RealPC: (a) ODGNet [6], (b) PointTr [62], (c) AdaPoinTr [62], (d) FoldingNet [60], (e) PCN [64], (f) TopNet [46], (g) SnowflakeNet [56], (h) GRNet [58], and (i) AnchorFormer [13]. These results are demonstrated in Table 4 and Figure 5.
In Table 3 (Right), we further benchmark RealPC on the task of reconstruction using standard LIDAR backbones: (a) DGCNN [52], (b) FoldingNet[61], and (c) the best performing completion method (Table 4), i.e. SnowflakeNet.
Finally, we benchmark RealPC on a shape generation strong diffusion-based model LION [50] in Table 3 (Left).
RealPC results on the completion task (Table 4 and Figure 5), and on generation and reconstruction tasks (Table 3) demonstrate the complexity of RealPC and the performance gap between it and existing object PC datasets.
Figure 5: Sub-optimal completion results on a RealPC PC using the best-performing completion baseline SnowflakeNet (Table 4).
@@ 4.4 Benchmarking on Neural Methods
We discuss the reasons behind the unsatisfactory performance of the baselines as follows: (a) We investigate existing PC datasets and RealPC using tools from algebraic topology and TDA in Section 4.2. Our discoveries, though striking, are not surprising. It suggests that RealPC consists of non-trivial 0- and 1-dimensional Homology-based topological features that correspond to connected components, and cycles - indicated by the non-diagonal nature of the persistence diagram in Figure 1. These add to the complexity of RealPC and provide strong evidence of the existence of higher dimensional topological features. We demonstrate that adding these Homology priors as constraints to an existing model can improve completion to some degree (Sec. 5.1 and Table 5). (b) PCs in RealPC are collected from four different sensors. This is a realistic setting where different incomplete PCs have been acquired with sensors with different intrinsic characteristics. Hence, different PC samples in such a case may follow different distributions, which is not true for existing datasets. (c) Section 4.2 shows that RealPC has several characteristics that make them challenging to work with. The point cloud acquisition in the real-world is affected by multiple ungovernable parameters. These factors introduce noise, non-uniform and inconsistent patterns - a characteristic that is absent in current object PC datasets.
5 Methodology
5.1 Persistent Homology Regularized Completion
Section 4.2 establishes that RealPC consists of rich topological features unlike existing datasets. We now focus on validating the use of 0-dim $\mathcal{PH}$ priors on top of an existing benchmarked method, to improve completion performance. We choose a PC completion model that can be easily integrated with $\mathcal{PH}$. There exist certain criteria that allow an existing model to be easily integrated with $\mathcal{PH}$. A challenge with $\mathcal{PH}$ is that it is compute-intensive. The creation of the Vietoris-Rips complex on a PC and persistence computation is time and compute intensive - number of simplices in the complex increase exponentially with increase in the PC size. Models that transform the input PC into seed PCs (at an intermediate model layer) at lower resolutions, generating multiple and sparse PC seeds, can allow diverse and compute-efficient calculation of $\mathcal{PH}$ priors. Such models are suited for integration of $\mathcal{PH}$ priors. We find ODGNet [6] to be a feasible model with multiple sparse seed point clouds at the decoder, which can be easily integrated with topological priors using moderate compute. The availability of these sparse seeds as learnable parameters at intermediate layers helps in training ODGNet using $\mathcal{PH}$-based topological priors.
5.1.1 Architecture
@@
We describe our methodology for the integration of ODGNet with topological priors - TopODGNet. We extract the multi-level seeds from the mid-level, low-level, and global features generated at the decoder. Seed point clouds consisting of points in the range of $256$ to $1024$ can be extracted at the decoder (Figure 6). The sparse nature of these seeds makes them easy to work with, and also allows efficient computation of topological features.
@@
Figure 6: TopODGNet. We calculate 0-dim $\mathcal{PH}$ based topological priors over sparse seeds and integrate it into the loss function. It enables completion along a topologically consistent skeleton.
@@ 5.1.2 Topological Loss
Figure 7: Left: A complete PC Middle: Visualization of the 00-dim $\mathcal{PH}$ based skeleton on a sparse seed during initial training. Right: 00-dim $\mathcal{PH}$ skeleton using sparse seeds as the training progresses. These priors assist the model in generating topologically consistent complete PCs.
@@
Given an incomplete PC, our goal is to augment it with new points that follow the global topology of the ground truth complete PC. A part of this work is handled by the dictionary module of ODGNet [6]. It ensures that the seed PCs are able to capture the artifacts of the complete PC and look similar to it. We extract the sparse multi-level PC seeds at the decoder (Figure 6). We now apply topological regularization priors on the seed PCs. To obtain this global topology-based backbone we use 0-dim $\mathcal{PH}$ priors on the seed PCs. 0-dim $\mathcal{PH}$ ensures the extraction of a complete PC skeleton (Fig. 7) that can serve as a prior which the network can follow, for complete PC generation. The underlying idea is to ensure shape consistency of the complete output along the skeleton, by generating points along the skeleton. We explain the process that extracts 0-dim $\mathcal{PH}$ priors from the seeds. The seed PC is converted into a simplicial complex. The initial complex is just the original PCs with an $\alpha$-radius ball (here $\alpha$=0). Increasing $\alpha$ adds edges, faces to the simplicial complex, thereby leading to the evolution and death of k-dim topological features. Each k-dim topological feature is characterized by a (birth, death) pair.
@@
Our goal is to use 0-dim $\mathcal{PH}$ based topological features. These focus on the global topology of the PC, providing a global skeleton that outlines the complete PC (Figure 7).
@@
Each 0-dim homology feature generated using filtration consists of a separate set of (birth, death) pairs denoted as b, d. The persistence of each pair is (b-d) (Section 3).
@@
We now explain our topological loss function. For 0-dim $\mathcal{PH}$, we minimize the sum of persistence for all extracted 0-dim persistence pairs. This ensures that at the end of the filtration, there is one connected component left (we minimize the persistence of all the pairs and there always exists at least one default component in the persistence diagram). This persistence diagram outlines the skeleton of the complete PC, which when used as a regularizer to a model, guides point generation along the skeleton (Figure 7). This is a vital observation - the network is able to observe sparse precise backbones of the complete PC during training which allow it to generate points more closer to the backbone. This observation is of immense interest and we utilize it in the next section to generate accurate complete PCs. Given $(b_{i},d_{i})$ refering to 0-dim $\mathcal{PH}$ based $(birth,death)$ pairs of a given seed PC, the topology loss is defined as follows:
$Topo\;Loss=\sum_{i=0}^{n}1\{i>k\}(b_{i}-d_{i})\\ =\sum_{i=k+1}^{n}(b_{i}-d_{i})$ (1)
@@
We explain the significance of $k$ now. The input partial PC may not necessarily have a single component i.e. it may be split into more than one component (Fig. 6). In such cases, $k\geq$ 2, which allows $\mathcal{PH}$ to generate multiple skeletons, each of which can attend to partial components in the input.
To establish that 0-dim $\mathcal{PH}$ priors demonstrate a similar effect on other real world datasets, we test these priors on an extensive scene-level dataset (KITTI) for scene completion and reconstruction (details in the supplementary material).
5.2 BOSH-Backbone Outline Sampler for PH
In Section 5.1, we studied the effect of 0-dim $\mathcal{PH}$ induced backbones as priors. A major shortcoming of 5.1.2 is the computational complexity of $\mathcal{PH}$. A single backbone requires generating the Vietoris-Rips complex of the PC (Figure 2). This is time and compute intensive as the number of simplices increase exponentially with the PC size. Further, the benefit of 5.1.2 is minor, as will be seen in Table 5 in Section 6 later.
0-dim backbone priors are proxies to ground truth shapes. We hypothesize that these can be directly sampled from the complete shapes’ surfaces at multiple sparsity levels. We circumvent the costly $\mathcal{PH}$ computations by introducing a simple and novel Homology Sampler (BOSH). BOSH directly samples numerous backbones from the surface of the ground truth PC (Figure 8). These act as proxies to the global 0-dim $\mathcal{PH}$ backbone. These sampled backbones also serve as seeds to a model, which grow into a complete shape PC as the training proceeds. Introduction of these seed backbones serves a dual purpose: (a) It enables the model to have a fine-grained overview of the complete shape at various resolutions. Given that sparse backbones are more difficult to complete, compared to the original shape, it forces the model to generate explicit attention to the precise shape details of the backbone for completion. (b) This strategy allows circumventing costly Vietoris-Rips complex computations.
Let $\{c_{i},p_{i}\}_{i=1}^{n}$ be the set of complete and incomplete PC pairs and $BOSH$ be the 0-dim Homology Sampler. M (here Chamfer Distance) refers to the similarity metric used, Net refers to the model and k refers to the number of Homology backbones sampled. The new total loss function is given by:
Figure 8: BOSHNet. Our compute-efficient Homology Sampler samples proxy $\mathcal{PH}$ backbones from the surface of the complete scan. These guide the completion process from the start of training.
$\small\begin{split}\sum_{i=1}^{n}\sum_{j=1}^{k}M(Net(BOSH(c_{i},j)),c_{i})+\sum_{i=1}^{n}M(Net(p_{i}),c_{i})\end{split}$ (2)
6 Experiments